# Heights and periods of algebraic cycles in families

### v1

@@ 6.3. Bigness of the generic fiber of the adelic line bundle
$\int_{{\mathscr{S}}_{i,q}\setminus\Omega_{i}}|c_{1}(\overline{\mathscr{L}}_{i})^{d}|\leq\int_{{\mathscr{S}}({\mathbb{C}})\setminus\Omega_{i}}c_{1}(\overline{\mathscr{E}}_{i}+\overline{\mathscr{F}}_{i})^{d}.$
Now let $\Omega_{i}^{\prime}\subset\Omega_{i}$ be another increasing sequence of relatively compact open subsets so that $\cup\Omega_{i}^{\prime}={\mathscr{S}}({\mathbb{C}})$. Then we can construct an increasing sequence of continuous functions $f_{i}$ so that $f_{i}(x)=1$ on ${\mathscr{S}}({\mathbb{C}})\setminus\Omega_{i}$ and $f_{i}=0$ on $\Omega_{i}^{\prime}$. Then for any $i\geq j$ we have
$\int_{{\mathscr{S}}({\mathbb{C}})\setminus\Omega_{i}}c_{1}(\overline{\mathscr{E}}_{i}+\overline{\mathscr{F}}_{i})^{d}\leq\int_{{\mathscr{S}}({\mathbb{C}})}f_{i}\cdot c_{1}(\overline{\mathscr{E}}_{i}+\overline{\mathscr{F}}_{i})^{d}\leq\int_{{\mathscr{S}}({\mathbb{C}})}f_{j}\cdot c_{1}(\overline{\mathscr{E}}_{i}+\overline{\mathscr{F}}_{i})^{d}.$
@@
$\limsup_{i{\longrightarrow}\infty}\int_{{\mathscr{S}}({\mathbb{C}})\setminus\Omega_{i}}c_{1}(\overline{\mathscr{E}}_{i}+\overline{\mathscr{F}}_{i})^{d}\leq\int_{{\mathscr{S}}({\mathbb{C}})}f_{j}\cdot c_{1}(\overline{\mathscr{E}}+\overline{\mathscr{F}})^{d}\leq\int_{{\mathscr{S}}({\mathbb{C}})\setminus\Omega_{j}^{\prime}}c_{1}(\overline{\mathscr{E}}+\overline{\mathscr{F}})^{d}.$
Let $j{\longrightarrow}\infty$. We get
$\lim_{i{\longrightarrow}\infty}\int_{{\mathscr{S}}({\mathbb{C}})\setminus\Omega_{i}}c_{1}(\overline{\mathscr{E}}_{i}+\overline{\mathscr{F}}_{i})^{d}=0.$
We are done. ∎
@@ A.2. Mumford–Tate group
It is clear that $h({\mathbb{S}})$ stabilizes this exact sequence and induced identity on $N$. Let $\mathbf{G}_{0}$ denote the Mumford–Tate group of $M_{-1}$. Then, we have an exact sequence of reductive groups:
$0{\longrightarrow}V{\longrightarrow}\mathbf{G}{\longrightarrow}\mathbf{G}_{0}{\longrightarrow}0,$
where $V$ is a vector group included into ${\mathrm{Hom}}(M_{0},M_{-1})$. If $M_{-1}$ is polarized, then $\mathbf{G}_{0}$ is reductive by Deligne. Thus, $V$ is the unipotent radical of $\mathbf{G}$.
The base change to ${\mathbb{R}}$ has a natural splitting $M_{\mathbb{R}}=M_{-1,{\mathbb{R}}}\oplus M_{0,{\mathbb{R}}}$ of Hodge structures given by the inverse of the isomorphism $M_{\mathbb{R}}\cap F^{0}M_{\mathbb{C}}{\overset{\sim}{{\longrightarrow}}}M_{0,{\mathbb{R}}}$. This induces splittings
$\mathbf{G}_{\mathbb{R}}=V_{\mathbb{R}}\rtimes\mathbf{G}_{0,{\mathbb{R}}}.$
@@
    The unipotent radical $V$ of $\mathbf{G}$ is a vector group.
@@ A.3. Variation of mixed Hodge structures and admissibility
    Any VMHS arising from geometry is admissible [EZ86] (see also [BZ14]).
@@ B.1. Classifying space
Let $M_{-1}$ be a finite dimensional ${\mathbb{Q}}$-vector space. Let $M=M_{-1}\oplus{\mathbb{Q}}$.
@@ B.1.1. Pure Hodge structures
Consider the polarized Hodge data on $M_{-1}$: a non-degenerate skew pairing $Q_{-1}\colon M_{-1}\otimes M_{-1}\rightarrow{\mathbb{Q}}(1)$, and a partition $\{h^{p,q}_{M_{-1}}\}_{p,q\in{\mathbb{Z}}}$ of $\dim M_{-1,{\mathbb{C}}}$ into non-negative integers with $p+q=-1$ such that $h^{p,q}_{M_{-1}}=h^{q,p}_{M_{-1}}$. Then there exists a classifying space ${\mathscr{M}}_{0}$ parametrizing ${\mathbb{Q}}$-Hodge structures on $M_{-1}$ of weight $-1$ with a polarization by $Q_{-1}$ such that the $(p,q)$-constituent of $M_{-1,{\mathbb{C}}}$ has complex dimension $h^{p,q}$. Moreover, the ${\mathbb{Q}}$-group $\mathbf{G}_{0}^{{\mathscr{M}}}:=\mathrm{Aut}(M_{-1},Q_{-1})$, the associated real Lie group $\mathbf{G}_{0}^{{\mathscr{M}}}({\mathbb{R}})^{+}$ acts transitively on ${\mathscr{M}}_{0}$, i.e.
@@
(B.1) $M_{-1,{\mathbb{C}}}=\bigoplus\nolimits_{p+q=-1}(M_{-1,x_{0}})^{p,q},\qquad F^{p}_{x_{0}}M_{-1,{\mathbb{C}}}=\bigoplus\nolimits_{p^{\prime}\geq p}(M_{-1,x_{0}})^{p^{\prime},q^{\prime}}$
with $(M_{-1,x_{0}})^{q,p}=\overline{(M_{-1,x_{0}})^{p,q}}$. The inclusion ${\mathscr{M}}_{0}\subseteq{\mathscr{M}}_{0}^{\vee}$ is given by $x_{0}\mapsto F^{\bullet}_{x_{0}}$.
For the Deligne torus ${\mathbb{S}}=\mathrm{Res}_{{\mathbb{C}}/{\mathbb{R}}}{\mathbb{G}}_{\mathrm{m},{\mathbb{C}}}$, the bi-grading decomposition above defines a morphism $h_{x_{0}}\colon{\mathbb{S}}\rightarrow\mathrm{GL}(M_{-1,{\mathbb{R}}})$, with $(M_{-1,x_{0}})^{p,q}$ the eigenspace of the character $z\mapsto z^{-p}\bar{z}^{-q}$ of ${\mathbb{S}}$. It is known that $h_{x_{0}}({\mathbb{S}})<\mathbf{G}_{0}^{{\mathscr{M}}}({\mathbb{R}})$ for all $x_{0}\in{\mathscr{M}}_{0}$. Hence we have a $\mathbf{G}_{0}^{{\mathscr{M}}}({\mathbb{R}})^{+}$-equivariant map, which is known to be injective,
@@
We close this subsubsection with the following remark on the Mumford–Tate group $\mathrm{MT}_{x_{0}}$ of the pure Hodge structure on $M_{-1}$ determined by $x_{0}$. We have that $\mathrm{MT}_{x_{0}}$ is a subgroup of $\mathbf{G}_{0}^{{\mathscr{M}}}$ for all $x_{0}\in{\mathscr{M}}_{0}$ and equals $\mathbf{G}_{0}^{{\mathscr{M}}}$ for some $x_{0}\in{\mathscr{M}}_{0}$.
@@ B.1.2. Weight $-1$ and 00
Fix the following data on $M=M_{-1}\oplus{\mathbb{Q}}$: the weight filtration $W_{\bullet}:=(0=W_{-2}M\subseteq W_{-1}M=M_{-1}\subseteq W_{0}M=M)$; the partition $\{h^{p,q}\}_{p,q\in{\mathbb{Z}}}$ of $\dim M_{{\mathbb{C}}}$ into non-negative integers, with $h^{p,q}=h^{p,q}_{M_{-1}}$ for $p+q=-1$ and $h^{0,0}=1$ and $h^{p,q}=0$ otherwise.
There exists the classifying space ${\mathscr{M}}$ parametrizing ${\mathbb{Q}}$-mixed Hodge structures $(M,W_{\bullet},F^{\bullet})$ of weight $-1$ and 00 such that:
@@
    the $(p,q)$-constituent $\mathrm{Gr}^{p}_{F}\mathrm{Gr}^{W}_{p+q}M_{\mathbb{C}}$ has complex dimension $h^{p,q}$;
@@
    $\mathrm{Gr}^{W}_{-1}M=M_{-1}$ is polarized by $Q_{-1}$.
See for example [Pea00, below (3.7) to the Remark below Lem. 3.9]. Notice that $\mathrm{Gr}^{W}_{0}M={\mathbb{Q}}$ is polarized by $Q_{0}\colon{\mathbb{Q}}\otimes{\mathbb{Q}}\rightarrow{\mathbb{Q}}$ is $a\otimes b\mapsto ab$. So $M$ is graded-polarized.
In our case, we need a better understanding of the structure of ${\mathscr{M}}$ than [Pea00]. The map $x\mapsto F_{x}^{\bullet}$ realizes ${\mathscr{M}}$ as a semi-algebraic open subset of a suitable flag variety ${\mathscr{M}}^{\vee}$, which is easily seen to be an orbit under $\mathbf{G}({\mathbb{C}})$ for the ${\mathbb{Q}}$-group
(B.3) $\mathbf{G}^{{\mathscr{M}}}:=M_{-1}\rtimes\mathrm{Aut}(M_{-1},Q_{-1})=M_{-1}\rtimes\mathbf{G}_{0}^{{\mathscr{M}}}.$
@@
As in the pure case, for the Mumford–Tate group $\mathrm{MT}_{x}$ for the mixed Hodge structure on $M$ determined by $x$, we have that $\mathrm{MT}_{x}<\mathbf{G}^{{\mathscr{M}}}$ for all $x\in{\mathscr{M}}$, and $\mathrm{MT}_{x}=\mathbf{G}^{{\mathscr{M}}}$ for some $x\in{\mathscr{M}}$.
@@
For the quotient $p\colon\mathbf{G}^{{\mathscr{M}}}\rightarrow\mathbf{G}_{0}^{{\mathscr{M}}}=\mathbf{G}^{{\mathscr{M}}}/M_{-1}$, consider the following surjective $\mathbf{G}^{{\mathscr{M}}}({\mathbb{R}})^{+}$-equivariant map
(B.6) $p\colon{\mathscr{M}}\rightarrow{\mathscr{M}}_{0},\quad g\cdot x\mapsto p(g)\cdot\bar{h}_{x}$
with $\bar{h}_{x}=p\circ h_{x}\colon{\mathbb{S}}\rightarrow\mathbf{G}^{{\mathscr{M}}}_{{\mathbb{R}}}\rightarrow\mathbf{G}^{{\mathscr{M}}}_{0,{\mathbb{R}}}$. This map sends the mixed Hodge structure $(M,W_{\bullet}M,F^{\bullet}M_{{\mathbb{C}}})$ to the associated pure Hodge structure on $M_{-1}=\mathrm{Gr}_{W}^{-1}M$.

### v2

@@ 6.3. Bigness of the generic fiber of the adelic line bundle
(6.11) $\int_{S_{i,q}\setminus\Omega_{i}}|c_{1}(\overline{\mathscr{L}}_{i,{\mathbb{C}}})^{\wedge d}|\leq\int_{S({\mathbb{C}})\setminus\Omega_{i}}c_{1}(\overline{\mathscr{E}}_{i,{\mathbb{C}}}+\overline{\mathscr{F}}_{i,{\mathbb{C}}})^{\wedge d}.$
Now let $\Omega_{i}^{\prime}\subseteq\Omega_{i}$ be another increasing sequence of relatively compact open subsets so that $\cup\Omega_{i}^{\prime}=S({\mathbb{C}})$. Then we can construct an increasing sequence of continuous functions $f_{i}$ so that $f_{i}\equiv 1$ on $S({\mathbb{C}})\setminus\Omega_{i}$ and $f_{i}\equiv 0$ on $\Omega_{i}^{\prime}$. Then for any $i\geq j$ we have
$\int_{S({\mathbb{C}})\setminus\Omega_{i}}c_{1}(\overline{\mathscr{E}}_{i,{\mathbb{C}}}+\overline{\mathscr{F}}_{i,{\mathbb{C}}})^{\wedge d}\leq\int_{S({\mathbb{C}})}f_{i}\cdot c_{1}(\overline{\mathscr{E}}_{i,{\mathbb{C}}}+\overline{\mathscr{F}}_{i,{\mathbb{C}}})^{\wedge d}\leq\int_{S({\mathbb{C}})}f_{j}\cdot c_{1}(\overline{\mathscr{E}}_{i,{\mathbb{C}}}+\overline{\mathscr{F}}_{i,{\mathbb{C}}})^{\wedge d}.$
@@
$\limsup_{i\rightarrow\infty}\int_{S({\mathbb{C}})\setminus\Omega_{i}}c_{1}(\overline{\mathscr{E}}_{i,{\mathbb{C}}}+\overline{\mathscr{F}}_{i,{\mathbb{C}}})^{\wedge d}\leq\int_{S({\mathbb{C}})}f_{j}\cdot c_{1}(\overline{\mathscr{E}}+\overline{\mathscr{F}})^{\wedge d}\leq\int_{S({\mathbb{C}})\setminus\Omega_{j}^{\prime}}c_{1}(\overline{\mathscr{E}}+\overline{\mathscr{F}})^{\wedge d}.$
Letting $j\rightarrow\infty$, we get
(6.12) $\lim_{i\rightarrow\infty}\int_{S({\mathbb{C}})\setminus\Omega_{i}}c_{1}(\overline{\mathscr{E}}_{i,{\mathbb{C}}}+\overline{\mathscr{F}}_{i,{\mathbb{C}}})^{\wedge d}=0.$
Now (6.9) follows immediately from (6.10), (6.11) and (6.12). We are done. ∎
@@ A.2. Mumford–Tate group
It is clear that $h({\mathbb{S}})$ stabilizes this exact sequence and induced identity on $M_{0}$. Let $\mathbf{G}_{0}$ denote the Mumford–Tate group of $M_{-1}$. Then, we have an exact sequence of reductive groups:
$0{\longrightarrow}\mathbf{G}_{-1}{\longrightarrow}\mathbf{G}{\longrightarrow}\mathbf{G}_{0}{\longrightarrow}0,$
where $\mathbf{G}_{-1}$ is a vector group included into ${\mathrm{Hom}}(M_{0},M_{-1})$. If $M_{-1}$ is polarized, then $\mathbf{G}_{0}$ is reductive by Deligne. Thus, $\mathbf{G}_{-1}$ is the unipotent radical of $\mathbf{G}$.
The base change to ${\mathbb{R}}$ has a natural splitting $M_{\mathbb{R}}=V_{{\mathbb{R}}}\oplus M_{0,{\mathbb{R}}}$ of Hodge structures given by the inverse of the isomorphism $M_{\mathbb{R}}\cap F^{0}M_{\mathbb{C}}{\overset{\sim}{{\longrightarrow}}}M_{0,{\mathbb{R}}}$. This induces splittings
$\mathbf{G}_{\mathbb{R}}=\mathbf{G}_{-1,{\mathbb{R}}}\rtimes\mathbf{G}_{0,{\mathbb{R}}}.$
@@
    The unipotent radical of $\mathbf{G}$ is a vector group.
@@ A.3. Variation of mixed Hodge structures and admissibility
    Any VMHS arising from geometry is admissible ([SZ85] and [Kas86]).
@@ B.1. Classifying space
Let $V$ be a finite-dimensional ${\mathbb{Q}}$-vector space. Let $E=V\oplus{\mathbb{Q}}$.
@@ B.1.1. Pure Hodge structures
Consider the polarized Hodge data on $V$: a non-degenerate skew pairing $Q_{-1}\colon V\otimes V\rightarrow{\mathbb{Q}}(1)$, and a partition $\{h^{p,q}_{V}\}_{p,q\in{\mathbb{Z}}}$ of $\dim V_{{\mathbb{C}}}$ into non-negative integers with $p+q=-1$ such that $h^{p,q}_{V}=h^{q,p}_{V}$. Then there exists a classifying space ${\mathscr{M}}_{0}$ parametrizing ${\mathbb{Q}}$-Hodge structures on $V$ of weight $-1$ with a polarization by $Q_{-1}$ such that the $(p,q)$-constituent of $V_{{\mathbb{C}}}$ has complex dimension $h^{p,q}$. Moreover, the ${\mathbb{Q}}$-group $\mathbf{G}_{0}^{{\mathscr{M}}}:=\mathrm{Aut}(V,Q_{-1})$, the associated real Lie group $\mathbf{G}_{0}^{{\mathscr{M}}}({\mathbb{R}})^{+}$ acts transitively on ${\mathscr{M}}_{0}$, i.e.
@@
(B.1) $V_{{\mathbb{C}}}=\bigoplus\nolimits_{p+q=-1}(V_{x_{0}})^{p,q},\qquad F^{p}_{x_{0}}V_{{\mathbb{C}}}=\bigoplus\nolimits_{p^{\prime}\geq p}(V_{x_{0}})^{p^{\prime},q^{\prime}}$
with $(V_{x_{0}})^{q,p}=\overline{(V_{x_{0}})^{p,q}}$. The inclusion ${\mathscr{M}}_{0}\subseteq{\mathscr{M}}_{0}^{\vee}$ is given by $x_{0}\mapsto F^{\bullet}_{x_{0}}$.
For the Deligne torus ${\mathbb{S}}=\mathrm{Res}_{{\mathbb{C}}/{\mathbb{R}}}{\mathbb{G}}_{\mathrm{m},{\mathbb{C}}}$, the bi-grading decomposition above defines a morphism $h_{x_{0}}\colon{\mathbb{S}}\rightarrow\mathrm{GL}(V_{{\mathbb{R}}})$, with $(V_{x_{0}})^{p,q}$ the eigenspace of the character $z\mapsto z^{-p}\bar{z}^{-q}$ of ${\mathbb{S}}$. It is known that $h_{x_{0}}({\mathbb{S}})<\mathbf{G}_{0}^{{\mathscr{M}}}({\mathbb{R}})$ for all $x_{0}\in{\mathscr{M}}_{0}$. Hence we have a $\mathbf{G}_{0}^{{\mathscr{M}}}({\mathbb{R}})^{+}$-equivariant map, which is known to be injective,
@@
We close this subsubsection with the following remark on the Mumford–Tate group $\mathrm{MT}_{x_{0}}$ of the pure Hodge structure on $V$ determined by $x_{0}$. We have that $\mathrm{MT}_{x_{0}}$ is a subgroup of $\mathbf{G}_{0}^{{\mathscr{M}}}$ for all $x_{0}\in{\mathscr{M}}_{0}$ and equals $\mathbf{G}_{0}^{{\mathscr{M}}}$ for some $x_{0}\in{\mathscr{M}}_{0}$. Moreover, ${\mathrm{MT}}_{x_{0}}$ is a reductive group for each $x_{0}\in{\mathscr{M}}_{0}$.
@@ B.1.2. Weight $-1$ and 00
Fix the following data on $E=V\oplus{\mathbb{Q}}$: the weight filtration
$W_{\bullet}:=(0=W_{-2}E\subseteq W_{-1}E=V\subseteq W_{0}E=E);$
the partition $\{h^{p,q}\}_{p,q\in{\mathbb{Z}}}$ of $\dim E_{{\mathbb{C}}}$ into non-negative integers, with $h^{p,q}=h^{p,q}_{V}$ for $p+q=-1$ and $h^{0,0}=1$ and $h^{p,q}=0$ otherwise.
There exists the classifying space ${\mathscr{M}}$ parametrizing ${\mathbb{Q}}$-mixed Hodge structures $(E,W_{\bullet},F^{\bullet})$ of weight $-1$ and 00 such that:
@@
    the $(p,q)$-constituent $\mathrm{Gr}^{p}_{F}\mathrm{Gr}^{W}_{p+q}E_{\mathbb{C}}$ has complex dimension $h^{p,q}$;
@@
    $\mathrm{Gr}^{W}_{-1}E=V$ is polarized by $Q_{-1}$.
See for example [Pea00, below (3.7) to the Remark below Lem. 3.9]. Notice that $\mathrm{Gr}^{W}_{0}E={\mathbb{Q}}$ is polarized by $Q_{0}\colon{\mathbb{Q}}\otimes{\mathbb{Q}}\rightarrow{\mathbb{Q}}$ is $a\otimes b\mapsto ab$. So $E$ is graded-polarized.
In our case, we need a better understanding of the structure of ${\mathscr{M}}$ than [Pea00]. The map $x\mapsto F_{x}^{\bullet}$ realizes ${\mathscr{M}}$ as a semi-algebraic open subset of a suitable flag variety ${\mathscr{M}}^{\vee}$, which is easily seen to be an orbit under $\mathbf{G}^{{\mathscr{M}}}({\mathbb{C}})$ for the ${\mathbb{Q}}$-group
(B.3) $\mathbf{G}^{{\mathscr{M}}}:={\mathrm{Aut}}(E,Q_{-1},W)=V\rtimes\mathrm{Aut}(V,Q_{-1})=V\rtimes\mathbf{G}_{0}^{{\mathscr{M}}}.$
@@
As in the pure case, for the Mumford–Tate group $\mathrm{MT}_{x}$ for the mixed Hodge structure on $E$ determined by $x$, we have that $\mathrm{MT}_{x}<\mathbf{G}^{{\mathscr{M}}}$ for all $x\in{\mathscr{M}}$, and $\mathrm{MT}_{x}=\mathbf{G}^{{\mathscr{M}}}$ for some $x\in{\mathscr{M}}$. Unlike the pure case, $\mathrm{MT}_{x}$ is not a reductive group in general.
@@
For each $x\in{\mathscr{M}}$, we have an induced Hodge structure $p(x)$ on $V$ and a unique vector $q(x)\in V_{\mathbb{R}}$ such that $(q(x),1)\in F^{0}_{x}E_{\mathbb{C}}$. Thus, we have a Hodge decomposition
$E_{\mathbb{C}}=\bigoplus_{p+q=-1,0}E_{x}^{p,q},\qquad E_{x}^{p,q}:=F_{x}^{p}E_{\mathbb{C}}\cap\overline{F_{x}^{q}E_{\mathbb{C}}}.$
The sum over $p+q=-1$ is the Hodge decomposition on $V$ and $E_{x}^{0,0}={\mathbb{C}}(q(x),1)$. This decomposition induces a bijection
(B.6) $(q,p)\colon{\mathscr{M}}{\overset{\sim}{{\longrightarrow}}}V_{\mathbb{R}}\times{\mathscr{M}}_{0},\qquad x\mapsto(q(x),p(x))$
where $p(x)$ is the induced Hodge structure on $V=W_{-1}E$. For the inverse map, for each $(v,x_{0})\in V_{\mathbb{R}}\times{\mathscr{M}}_{0}$ we take $x\in{\mathscr{M}}$ be the unique Hodge structure $h$ on $E$ extending $x_{0}$ on $V_{\mathbb{C}}$ so that $F^{0}_{x}E_{\mathbb{C}}=F^{0}_{x_{0}}V_{\mathbb{C}}+{\mathbb{C}}(v,1)$. The real semi-algebraic structure on ${\mathscr{M}}$ is given by the product in the above isomorphism. The complex structure on ${\mathscr{M}}$ is given by making ${\mathscr{M}}$ into a vector bundle over ${\mathscr{M}}_{0}$ with fibers $V_{\mathbb{R}}{\overset{\sim}{{\longrightarrow}}}V_{\mathbb{C}}/F^{0}_{x}V_{\mathbb{C}}$. Notice that for each $a\in V({\mathbb{R}})$, the submanifold $(q,p)^{-1}(\{a\}\times{\mathscr{M}}_{0})$ is a semi-algebraic and complex submanifold of ${\mathscr{M}}$. By abuse of notation, we also denote the group quotient still by $p\colon\mathbf{G}^{{\mathscr{M}}}\rightarrow\mathbf{G}_{0}^{{\mathscr{M}}}=\mathbf{G}^{{\mathscr{M}}}/V$.
For each $x\in{\mathscr{M}}$, we have a surjective morphism ${\mathrm{MT}}_{x}{\longrightarrow}{\mathrm{MT}}_{p(x)}$ with kernel $V_{x}:=V\cap{\mathrm{MT}}_{x}$. Thus $V_{x}$ is the unipotent radical of ${\mathrm{MT}}_{x}$. This surjection has a section $s\colon{\mathrm{MT}}_{p(x)}{\longrightarrow}{\mathrm{MT}}_{x}$. Via (B.3), this section can be written as $s(g)=(\sigma(g),g)$ with $\sigma$ a cocycle in $Z^{1}({\mathrm{MT}}_{x},V)$. But $H^{1}({\mathrm{MT}}_{p(x)},V)=0$ since ${\mathrm{MT}}_{p(x)}$ is reductive. Thus $\sigma$ is a coboundary, i.e. $\sigma(g)=v_{0}-gv_{0}$ for some $v_{0}\in V$ unique up to addition by invariant vectors of $V$ under action by ${\mathrm{MT}}_{p(x)}$. Thus, we get a precise Levi decomposition
(B.7) ${\mathrm{MT}}_{x}=V_{x}\rtimes{\mathrm{MT}}_{x_{0}}={\mathrm{Ad}}(v_{0})(V_{x}\rtimes{\mathrm{MT}}_{p(x)}),$
where $V_{x}=V\cap{\mathrm{MT}}_{x}$, $x_{0}=(v_{0},p(x))$ is a point of ${\mathscr{M}}$ via the isomorphism (B.6). With this isomorphism, we see that $x=(v_{1}+v_{0},p(x))$ with $v_{1}\in V_{\mathbb{R}}$ which has Zariski closure $V_{x}$ over ${\mathbb{Q}}$.